40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 11 22:11:15 2015, Koji, Update, SUS, PSL/IOO maintenance, TM SUS check up TM_SUS.pdfTM_SUS_SD.pdf
    Reply  Thu Nov 12 09:15:56 2015, Steve, Update, SUS, ITMX osems ITMX.png
       Reply  Fri Nov 13 15:20:24 2015, Steve, Update, SUS, OSEMs OSEMs.pngPOSbiasOFF.png
    Reply  Thu Nov 12 09:41:02 2015, Steve, Update, PSL, PMC PMC_ITMX.png
    Reply  Tue Nov 17 00:57:21 2015, ericq, Update, SUS, ITMX UL calmed? TM_nov17.pdfITMX_sensortrend.png
       Reply  Tue Nov 17 15:59:23 2015, rana, Update, SUS, ITMX UL calmed? 
Message ID: 11760     Entry time: Fri Nov 13 15:20:24 2015     In reply to: 11755
Author: Steve 
Type: Update 
Category: SUS 
Subject: OSEMs 

Are they oscillating or not ?

Quote:

The ITMX OSEMS are oscillating.

ETMX, ETMY and MC2 POS _biases are off. Why ?

The Epics MEDMs screens are going blank in  ~3-5 minutes reluctantly.

Attachment 1: OSEMs.png  467 kB  | Hide | Hide all
OSEMs.png
Attachment 2: POSbiasOFF.png  87 kB  | Hide | Hide all
POSbiasOFF.png
ELOG V3.1.3-